OpenStudy (brunsonni):

Solve 2x2 + 5x + 5 = 0. Round solutions to the nearest hundredth. No real solutions x ≈ -3.27 and x ≈ 0.77 x ≈ -2.98 and x ≈ -7.02 x ≈ -0.77 and x ≈ 3.27

OpenStudy (brunsonni):

2x2 is 2x^2

OpenStudy (campbell_st):

you can use the determinant... to check the solutions or use the general quadratic formula to solve this

OpenStudy (brunsonni):

so a=2, b=5, and c=5, yeah?

OpenStudy (brunsonni):

@campbell_st is it A, because i got x=-0.43, and x=-2.07 and those arent options.

OpenStudy (campbell_st):

well you shouldn't have been able to get answers... as the discriminant in this question is \[b^2 - 4ac = 5^2 - 4 \times 2 \times 5 = -15\] and you can't take the square root of a negative number... so you choice of A is correct. but your method is flawed

OpenStudy (brunsonni):

thanks